OpenStudy (anonymous):

sqrt{b2-4ac};b=3,a=-1,c=-2

Parth (parthkohli):

You just plug in values, so if you are given a question like 73x^2;x=2 then you will plug the 2 inside x. 73(2)^2

OpenStudy (lgbasallote):

substitute \[\sqrt{(3)^2 - 4(-1)(-2)} \implies \sqrt{9 - 8}\]

Parth (parthkohli):

So, you plug in and get \( \sqrt{3^2 - 4(-1)(-2)}\)

Parth (parthkohli):

@rannsan Why did you close that question?